# Survey on Modeling of Articulated Objects

### v1

@@ 4.2.1 Intermediate Representation
  • Image feature (image feat.). Image latent features are used as the intermediate representation in the works that take RGB or RGB-D images as input. Similar to the point cloud features, the image features are extracted from pixels for later use in mobility part analysis. Some works follow this scheme to extract 2D features from depth images instead of per-point features from point clouds. ResNet [HZRS16] is the common backbone used for the feature extraction in the existing works.
Articulated object creation. For the work focusing on articulated object creation, the neural field is the most commonly used intermediate representation. A neural field is an implicit function parameterized by a neural network to represent the 3D geometry of the object in a continuous space. The network itself can be regarded as an SDF, occupancy field, neural density field, or neural radiance field, depending on the way the network is trained. Once constructed, the neural field can be queried to reconstruct or generate the objects and further be extracted to the mesh representation. To be more specific, A-SDF [MQK*21], [WCM*22, ], CARTO [HIZ*23] leverage neural SDF to represent the 3D geometry of the object. Ditto [JHZ22] constructs a neural occupancy field to indicate the occupancy status for each articulated part. CLA-NeRF [TLYS22] and PARIS [LMS23] both use a neural radiance field to represent both the shape and appearance of each part of the object.
@@ 4.2.2 Strategy and Supervision Signal
In this section, we discuss the different choices of algorithms or learning strategies adopted by the works when analyzing the articulated part structure.
Handcrafted methods. To infer part structure from the input geometry, early works rely on heuristic rules, human-designed descriptors, or mathematical algorithms to analyze the shape.
For the works that take a single-state mesh as input, the approaches based on the slippage analysis, geometric properties, and relationship descriptors are adopted to effectively segment the parts associated with the articulated motion. The slippage analysis is proposed by [GG04, ] to analyze the shape by discovering the slippable motions. It can segment the geometry into slippable portions which they call kinematic surfaces to associate with rigid motions for each part. Leveraging this approach, [XWY*09, ] detected primitive surfaces from the input object and link to motion joints that connect the articulated parts. Geometric attributes on the mesh surfaces can be characteristic for inferring different joint configurations. Based on this insight, [MYY*10, ] proposed to characterize the part motion by detecting the self-similarity and symmetry on the mesh surface for each mechanical part in a mathematical way. Spatial relationships between the parts are also informative for segmenting the components. Building on this observation. [SHL*14, ] proposed to leverage the supporting-supported relationship between surface patches to decompose a scene into a set of objects and their articulated parts. Similarly, [HLV*17, ] proposed to leverage the interaction pattern between parts described by a human-designed interaction descriptor [HZvK*15]. Based on these geometric features, a distance function is designed to measure the similarity between snapshots of the object.
For the works that take point clouds observing a motion sequence as input, mathematical algorithms, and handcrafted descriptors are the two common choices to find the correspondence between adjacent frames. Once the correspondence is established, the part segmentation can be inferred from the 3D trajectory by clustering the points that are moving together. Given a known skeleton of the object in the first frame, [PG08, ] proposed to trace the points that are skinned on each bone in the other frames using the Iterated Closest Point (ICP) algorithm [BM92]. Consuming an unlabelled point cloud sequence, [YLX*16, ] and [SCZ21, ] proposed to leverage a deformable 3D shape registration algorithm [PB11] to estimate the 3D trajectory. Taking RGB-D sequence with additional color information as input, [LWL*16, ] proposed to use scene flow [JSGC15] to produce a raw dense correspondence, from which SIFT features [Low04] are extracted from RGB images to refine and prune the initial proposals in the dense trajectory.
@@
Figure 6: Shape2Motion [WZS*19] method pipeline that jointly segments the articulated parts and estimates motion attributes from a single point cloud as input. Figure reproduced from original paper [WZS*19].
@@
Supervised learning methods. For the task of articulated part perception, most of the existing works choose to learn the part assignment from a single state of the object in a supervised manner. Under this setting, the part segmentation is essentially formulated as a classification problem, where the goal is to predict the part label for each point in the input shape by transferring the knowledge learned from training data. Most of the work under this group takes the point cloud optionally with color as input, and the classification model is trained with supervision from signals in 3D that are available from the human-annotated datasets. Shape2Motion [WZS*19] is the pioneering work that contributes both the dataset and the learning model for the task. Their network is trained to propose motion parts and attributes followed by a matching module to refine the final prediction. Figure 6 illustrates the their method pipeline. With more 3D data with part annotations available, more works [YHY*19, ATK19, LWY*20, AFA*22, LXX*22, GXZ*23] follow the trend of joint prediction for segmentation and motion with varied network designs. Note that the part semantics are not always aligned with the mobility, but the data with semantic segmentation labels can be an initial source to refine the articulated part segmentation. Inspired by this idea, [LSH*23, ] contribute to extending a semantic segmentation dataset PartNet [MZC*19] with more articulation annotation by transferring the knowledge trained on PartNet-Mobility [XQM*20] dataset in a supervised way. To make the segmentation robust to arbitrary articulated states of the object, a common practice is to augment the training objects with different articulation states. It requires the data to be annotated with the articulation parameters, which is expensive and not at scale. In an effort to reduce reliance on articulation annotation available, Banana [DLS*23] design a network to be aware of the inter-part equivariance so that the knowledge acquired from objects only in resting states can be generalized to any other articulated states. This strategy allows the datasets compiled from static objects with segmentation labels to be usable for articulated part perception. Another line of work in this category attempts to perceive the part structure from RGB(-D) images. Given a pair of segmented parts in a single-view RGB-D image, [ZLLK21, ] proposed FormNet to estimate connectedness and motion flow between the parts in a supervised way. OPD [JMSC22] and OPDMulti [SJSC23] proposed to learn 2D segmentations from single-view RGB images in a supervised manner, where the ground truth masks are available from the datasets they collected. [QJR*22, ] first proposed to infer mobility from an RGB video observing the human-object interaction. They train a network to regress 3D plane parameters to represent the articulated parts by leveraging the supervision from the 3D datasets without articulation annotation.
@@
For the task of articulated object creation, supervised learning methods are also widely adopted. A-SDF [MQK*21], [WCM*22, ], and CARTO [HIZ*23] is a line of work for the object-level reconstruction of a shape that is deformable to a specific state. The model training is supervised by either ground truth SDF or multi-view RGB images. These supervision signals guide the model to learn an underlying distribution of the geometry that is associated with the articulation states. To further achieve part-level reconstruction, Ditto [JHZ22] and CLA-NeRF [TLYS22] require additional supervision from the part segmentation labels either on the point cloud or the multi-view images.
Figure 7: Generative models for articulated object synthesis. NAP [LDS*23] unconditionally synthesizes articulated objects with a full description, and CAGE [LTMS23] learns to generate objects with part structure conforming to the conditional input graph. Figures reproduced from original papers [LDS*23, LTMS23].
@@
Self-supervised learning methods. Since the articulated part segmentation and articulation annotations are expensive to obtain, self-supervised learning methods have been proposed recently to reduce the reliance on these supervision signals. Taking a point cloud with a known number of parts as input, [LZH*23, ] proposed to learn the segmentation by extracting part-level equivariant features and factorizing the shape into canonical states. Through the iterative processing of canonicalization and then reconstruction using the factorized parameters, part decomposition can be learned by minimizing the reconstruction error on the input point cloud. PARIS [LMS23] reconstructs separate parts from a pair of multi-view RGB images that observe the object in two different states by leveraging differential rendering. By identifying the static and moving components through iterative optimization, the part-level geometry can be reconstructed by minimizing the photometric error between the input images and the rendered ones from the reconstructed surface.
Generative models. Leveraging the recent advance of denoising diffusion models [HJA20] in 3D generation, a line of work NAP [LDS*23] and CAGE [LTMS23] train generative models to synthesize the articulated object with part structure in a compositional way. The process of data distribution learning requires extensive data with complete annotations for articulated objects, such as a kinematic tree, part geometry represented as SDF in NAP [LDS*23] or as bounding primitives in CAGE [LTMS23]. See Figure 7 for the overview of the NAP and CAGE work. The acquisition of high-quality synthetic and real data with these annotations is still expensive and not at scale, which limits the generation ability of these models beyond the training data distribution.

### v3

@@ 4.2.1 Intermediate Representation
  • Image feature (image feat.). Image latent features are used as the intermediate representation in the works that take RGB or RGB-D images as input. Similar to the point cloud features, the image features are extracted from pixels for later use in mobility part analysis. ScrewNet jain2021screwnet follow this scheme to extract 2D features from depth images using ResNet he2016deep instead of per-point features from point clouds. For the work focused on articulation perception from a single image jiang2022opd; sun2023opdmulti; wang2025active, MaskRCNN he2017mask and Mask2Former cheng2022masked are the common backbones used to extract feature maps for detecting and masking the parts.
  • MV RGBs. Articulate AnyMesh qiu2025anymesh segments the input mesh by using muli-view images as proxy. Once the articulated parts are segmented in the rendered multi-view images, the part labels are projected back to the mesh by voting from the corresponding pixels. It introduces a effective way to leverage the advanced image segmentation models to assist the 3D segmentation task.
Articulated object reconstruction. In the context of articulated object reconstruction, neural fields have emerged as the predominant intermediate representation. A neural field is an implicit function parameterized by a neural network, enabling the continuous and differentiable representation of an object’s 3D geometry, with the optional capability to model its appearance. Once constructed, the neural field can be queried to reconstruct the objects and further converted into a mesh representation using the marching cube algorithm lorensen1998marching. Depending on its formulation, existing works represent the object or their parts with signed distance functions (SDFs), occupancy fields, neural radiance fields (NeRFs), or 3D Gaussians.
  • Neural SDF. SDFs define the geometry of an object by encoding the signed distance from any point in space to the object’s surface, with positive values outside the object and negative values inside. A-SDF mu2021asdf, wei2022self, CARTO heppert2023carto are the works that leverage neural SDF to represent the object as a singular surface. The compactness of SDFs makes them effective at capturing smooth surfaces and can be trained to be deformable to animate the object into different articulation states. However, it becomes challenging to represent multiple parts with discrete surfaces in a single SDF, which limits the ability of this line of work to model articulated objects with multiple parts and complex structures.
  • Neural occupancy field. Occupancy fields are binary functions that indicate whether a point in space is occupied by the object or not. Ditto jiang2022ditto constructs a neural occupancy field from a pair of point clouds to represent the object with additional labels for each point in the field to indicate the part association. Real2Code mandi2024real2code also learns an occupancy-based shape completion network to model part surfaces. Both methods use occupancy fields to represent articulated objects from point cloud inputs. Occupancy fields provide flexible representations which effectively handle non-watertight and fragmented structures. However, like SDFs, occupancy fields also require dense 3D sampling as supervision to capture the surface details, leading to high training costs.
  • NeRF. NeRF is a continuous volumetric representation that models the object’s geometry and appearance as a function of 3D coordinates. It is an effective representation that bridges the gap between 2D images and 3D representations via differentiable rendering. Several work uses NeRFs for articulated object reconstruct from multi-view RGB images tseng2022cla; liu2023paris; swaminathan2024leia, RGB-D images weng2024neural or RGB videos song2024reacto. The geometry of each part can be reconstructed by either learning separate NeRFs for each part liu2023paris or by associating part labels with the 3D coordinates tseng2022cla; weng2024neural; song2024reacto. LEIA swaminathan2024leia learns the interpolated state of the object between two given states by training a single NeRF. NeRFs excel at synthesizing realistic novel views of the objects under different articulated states once trained. But the quality of the extracted surface from NeRFs is easily affected by the sparsity of the training data, textureless regions, occlusions, and lighting conditions. So they often require densely captured views for accurate reconstruction.
  • 3D Gaussians. 3D Gaussian representation models a scene or object using a set of anisotropic Gaussian functions in 3D space, each defined by position, shape, opacity, and color. 3D Gaussian Splatting is a cutting-edge technique in neural rendering that enables efficient rendering, smooth interpolation, easy dynamic modeling, and high-quality reconstruction of the object. RSRD kerr2024robot first leverage this representation to model the dynamics of the articulated object over time as 4D Gaussian fields, which visually imitate the articulation from an object scan in multi-view and a monocular video. From the multi-view scan, they first reconstruct the static object using 3D Gaussians and cluster the Gaussians into semantic components. Each component is then tracked over time to capture the dynamics of the object parts by referring to the monocular video in DINO caron2021emerging feature space. ArtGS liu2025building leverages 3D Gaussians to reconstruct multi-part articulated objects given two states of the object in multi-view RGBD images. Compared to other more compact representations, such as SDFs, 3D Gaussians provide flexibility in representing multiple components by associating specific Gaussians with each part, enabling a more explicit way of capturing dynamics and interactions. 3D Gaussian representation presents a promising step toward faster and more efficient reconstruction of articulated objects, which worth further exploration in the future.
Figure 7: Selected examples of the intermediate representation used for articulated part perception and creation. Figures reproduced from the original papers mitra2010illustrating; li2016mobility; mescheder2019occnet; mildenhall2021nerf; kerbl20233d.
The advantage of using implicit fields lies in their continuous and differentiable properties, making them well-suited for applications such as inverse rendering, shape fitting and completion, and deformation modeling. However, a major drawback is their computational expense, as each scene or object requires overfitting or optimization on a separate network, which limits their scalability and efficiency for large-scale or real-time use cases. Real2Code mandi2024real2code bridges the 2D and 3D gap with depth estimation model wang2024dust3r and extract point cloud from the depth map. The point cloud are then labeled as different parts by projecting the segmentation from multi-view images by leveraging SAM kirillov2023sam, which then completes the part surfaces by training a shape completion network. This pipeline shows its advantage in computational efficiency and effectiveness by taking advantage of recent advances in foundational models. But this approach relies on minimal occlusion in input images, otherwise the part geometry would be largely incomplete.
Articulated object generation. In articulated object generation, designing an effective intermediate representation is crucial for capturing the complex structure and articulation of objects. Given their inherently structured nature, articulated objects can be recorded in formats like URDF or MJCF for simulation in physics engines. Inspired by this, the intermediate representation for the generation task is designed to encode the part configuration in a hierarchical manner. A common approach uses the part bounding box to anchor the spatial location of each part, while the surface geometry is compressed into a latent feature space if it is a part of the generation target. See Figure 6 for a visualization of an example representation proposed in NAP lei2023nap.
NAP lei2023nap pioneers the design of the hierarchical representation as mentioned above. To model the part surface, they trained an occupancy shape autoencoder mescheder2019occupancy to encode each part surface as a high-dimensional latent code which can be later decoded as SDFs. CAGE liu2024cage and SINGAPO liu2024singapo represent the parts with only bounding boxes with semantic labels as embeddings during the generation process, and then retrieve the part surface under the context of the generated part layout. This abstraction strategy allows the model to focus more on the structural arrangement and their interplay with the articulation at the generation stage. ArtFormer su2024artformer further extends the representation from NAP and CAGE. It proposes to encode the part point cloud into triplane features and quantize the features into a discrete codebook, which is then used as latent code to generate the part geometry along with the structure. The generated latent code can be decoded into SDFs to represent the part surface. MeshArt follows the hierarchical representation design but tailors the geometry representation to directly generate the mesh vertices from the latent code. As illustrated in  Figure 12, in the first generation stage, it represents each part’s bounding box as a triangle mesh and learns a codebook as quantized triangle embeddings. Guided by the global object structure generated in the first stage, another geometry codebook is learned for encoding the mesh triangles and their likelihood of being junction faces. This design allows the model to directly generate triangle meshes for articulated parts, resulting in sharp and clean output surfaces.
In contrast to the above works that model geometry and structure based on prior learning from training data distributions, Articulate-Anything le2024articulate proposes to represent objects as programs that configure the part composition and articulation parameters. An actor-critic system is developed to synthesize Python code that can be compiled into URDF files. By rendering the object synthesized with the programs, the vision-language critic can provide feedback to refine the program generation. This approach relies on the programmatic representation to ensure the plausibility of the generated object, which can be more interpretable and interactable than the learned latent space. However, like CAGE and SINGAPO, it relies on a 3D assets library to retrieve the part geometry, which may limit the diversity of the generated objects.
On the image-driven generation side, DragAPart li2024dragapart generates the part-level interaction in image space by encoding the drag-driven animation to the image features. They teach the pre-trained generative model to respond to drag prompt while enforcing physical consistency with respect to the underlying 3D geometry by fine-tuning the model with paired 2D images with drag annotations. The motion prior learned by the model can be used to generate the object images in different states by dragging the part in the image.
Discussion on intermediate representations. The choice of an intermediate representation depends on the input data modality, the application scenario, and the specific requirements of each task. For perception tasks, it is crucial to capture local geometric features and spatio-temporal relationships that enable part decomposition and motion analysis. Early work often relied on explicit forms (e.g., surface patches or 3D trajectories), while newer approaches employ implicit latent features from point clouds or images, reflecting a shift from handcrafted to learning-based techniques. For creation tasks, representations must balance interpretability, computational demands, fidelity, and ease of real-world data acquisition. Implicit neural fields are popular for their continuous, differentiable nature—valuable in tasks like inverse rendering—but they typically require dense supervision and struggle to represent discrete part structures. A promising direction is to blend discrete and continuous representations (e.g., 3D Gaussians), which allows modeling piecewise-rigid parts in a differentiable manner. The combination with the programmatic representation can further enhance the interpretability and controllability of the generated objects, which is beneficial for applications that require human interaction or physical simulation. These hybrid representations also lend themselves to hierarchical structures, which is especially useful for modeling more structurally complex objects with multiple parts in a kinematic chain. Along this direction, integrating graph representations or programs as domain-specific languages can be a promising approach to enhance the interpretability and controllability of generated objects with a hierarchical structure.
@@ 4.2.2 Strategy and Supervision Signal
In this section, we discuss different algorithms and learning strategies adopted by the works analyzing the articulated part structure.
Handcrafted methods. To infer part structure from the input geometry, early works rely on heuristic rules, human-designed descriptors, or geometric properties to analyze the shape.
For the works that take a single-state mesh as input, the approaches based on the slippage analysis, geometric properties, and relationship descriptors are adopted to effectively segment the parts associated with the articulated motion. The slippage analysis is proposed by gelfand2004shape to analyze the shape by discovering the slippable motions. It can segment the geometry into slippable portions which they call kinematic surfaces to associate with rigid motions for each part. Leveraging this approach, xu2009joint detected primitive surfaces from the input object and link to motion joints that connect the articulated parts. Geometric attributes on the mesh surfaces can provide valuable signals for inferring different joint configurations. Based on this insight, mitra2010illustrating proposed to characterize the part that related to motions by detecting the geometrical primitives with self-similarity and symmetry on the mesh surface for each mechanical part. Spatial relationships between the parts are also informative for segmenting the components. Building on this observation. sharf2014mobility proposed to leverage the supporting-supported relationship between surface patches to decompose a scene into a set of objects and their articulated parts. Similarly, hu2017learning proposed to leverage the interaction pattern between parts described by a human-designed interaction descriptor hu2015interaction. Based on these geometric features, a distance function is designed to measure the similarity between snapshots of the object.
For the works that take point clouds observing a motion sequence as input, geometric alignment algorithms, and handcrafted descriptors are the two common choices to find the correspondence between adjacent frames. Once the correspondence is established, the movement of each point over time can be tracked to construct the 3D trajectory. Then the part segmentation can be inferred from the 3D trajectory by clustering the points that are moving together. Given a known skeleton of the object in the first frame, pekelny2008articulated proposed to trace the points that are skinned on each bone in the other frames using the Iterative Closest Point (ICP) algorithm besl1992method. Consuming an unlabelled point cloud sequence, yuan2016space and shi2021self proposed to leverage a deformable 3D shape registration algorithm papazov2011deformable to estimate the 3D trajectory. Taking RGB-D sequence with additional color information as input, li2016mobility proposed to use scene flow jaimez2015primal to produce a raw dense correspondence, from which SIFT features lowe2004distinctive are extracted from RGB images to refine and prune the initial proposals in the dense trajectory.
@@
Figure 8: Shape2Motion wang2019shape2motion jointly segments the articulated parts and estimates motion attributes from a single point cloud as input. Figure reproduced from original paper wang2019shape2motion.
Supervised learning methods. On the perception side, a line of works proposes to learn the part decomposition from a single state of the object in a supervised manner. A key distinction between articulated part segmentation and semantic part segmentation is that an articulated part may not be semantically atomic and can sometimes be composed of multiple semantic components. Unlike semantic segmentation, where labels can be consistently assigned across object instances, defining generalizable labels for articulated parts across categories is more challenging. Shape2Motion wang2019shape2motion is the pioneering work that addresses the problem with a proposal-based solution, as illustrated in Figure 8. The model learns to regress a similarity matrix between points and a confidence score for each part proposal. Using ground-truth part grouping, the similarity is supervised to minimize the point-wise feature distance among the points that belong to the same part, while the confidence map is trained to maximize the IoU between the point set of the proposal and its ground truth counterpart. Building on this idea later works adopt similar feature grouping strategies but explore different network architectures to improve segmentation accuracy and efficiency, such as the recurrent-based network used in RPMNet yan2019rpm in capture temporal relationships and graph convolutional network used in abdul2022learning to learn the hierarchical part structure. While these methods demonstrate effective articulated part prediction across object categories, they heavily rely on large-scale labeled point cloud datasets for training.
Another line of work focuses on pose estimation for the articulated parts from a single state of the object, where part segmentation is a necessary first step. Early works make the assumption that the object category is known and all the instances in the category share the same articulated structure. Under this assumption, li2020category proposed to represent each class of objects in an articulation-aware normalized coordinate space hierarchy (ANCSH) (as shown in Figure 9). The network is trained to map the new object instance to the normalized coordinate space with a branch for assigning the part labels, which are supervised by the per-point ground truth labels. This method inspired a series of works weng2021captra; liu2022toward that leverage this idea to address the articulated part segmentation and pose estimation on the category level. The main limitation of these methods is that the assumption of shared part structure across instances in the same category may not hold for all object categories, which limits the generalization capability of the model. To address this limitation, geng2023gapartnet proposed to address the same problem in a category-agnostic manner by introducing generalizable actionable parts (GAPart). Using these ground-truth GAPart labels as supervision, the trained part segmentation model is more generalizable to unseen object categories.
Figure 9: Articulation-aware Normalized Coordinate Space Hierarchy (ANCSH) proposed by li2020category. ANCSH is a category-level representation that defines canonicalized object state with joint parameters in NAOCS, and normalizes each part in NPCS while maintaining the part orientation in NAOCS. The colors represent the corresponding coordinates in each space. Figure reproduced from original paper li2020category.
@@
In an effort to reduce reliance on articulation annotation available, Banana deng2023banana designs a network to be aware of the inter-part equivariance so that the knowledge acquired from objects only in resting states can be generalized to any other articulated states. This strategy allows the datasets compiled from static objects with segmentation labels to be usable for articulated part perception. Another line of work in this category attempts to perceive the part structure from RGB(-D) images. Given a pair of segmented parts in a single-view RGB-D image, zeng2021visual proposed FormNet to estimate connectedness and motion flow between the parts in a supervised way. OPD jiang2022opd, OPDMulti sun2023opdmulti and wang2025active propose to learn 2D segmentations from single-view RGB images in a supervised manner, where the ground truth masks are available from the datasets they collected. qian2022understanding first proposed to infer mobility from an RGB video observing the human-object interaction. They train a network to regress 3D plane parameters to represent the articulated parts by leveraging the supervision from the 3D datasets without articulation annotation.
@@
For the task of articulated object creation, supervised learning methods are also widely adopted. A-SDF mu2021asdf, wei2022self, and CARTO heppert2023carto propose a line of work for the object-level reconstruction of a shape that is deformable to a specific state. The model training is supervised by either ground truth SDF or multi-view RGB images. These supervision signals guide the model to learn an underlying distribution of the geometry that is associated with the articulation states. To further achieve part-level reconstruction, Ditto jiang2022ditto requires additional supervision from the part segmentation labels on point clouds, and CLA-NeRF tseng2022cla and Real2Code mandi2024real2code learn from segmentation masks on multi-view images.
Semi-weakly supervised learning methods. While part semantics do not always align with mobility, semantic segmentation labels can serve as a valuable source for refining articulated part segmentation. Inspired by this idea, liu2023semi contributes a semi-weakly supervised approach to tackle the task of object kinematic motion prediction problem. Their method leverages the fine-grained and hierarchical part labels from PartNet mo2019partnet as weak signals to train a graph neural network. It also prunes the input semantic hierarchy and extracts a mobility tree to represent the object’s articulated structure. Then parts predicted to have motion in the output mobility tree are used as pseudo-labels to expand the training data in the PartNet-Mobility xiang2020sapien dataset for a part mobility prediction network in a semi-supervised manner. This semi-weakly supervised pipeline demonstrates its advantage by improving model performance through the use of semantic annotations, which are generally more accessible than motion-part annotations. However, its effectiveness is limited by the granularity of the semantic hierarchy in the source dataset, as a coarse hierarchy may lead to imprecise mobility predictions. Additionally, the method relies on a sufficient amount of ground-truth labeled data to train the model before generating pseudo-labels for each category, as illustrated in Figure 10. The reliability of these pseudo-labels depends on the quality of the motion prediction network, and errors in motion estimation can introduce noise into the training process, potentially affecting overall model performance.
Figure 10: liu2023semi introduces a semi-weakly supervised learning approach that leverages the fine-grained and hierarchical part labels as weak signals to predict mobility trees, which are then used as pseudo-labels to train mobility analysis models. Figure reproduced from original paper liu2023semi.
@@
Self-supervised learning methods. Since the articulated part segmentation and articulation annotations are expensive to obtain, self-supervised learning methods have been proposed recently to reduce the reliance on these supervision signals. Taking a point cloud with a known number of parts as input, liu2023self proposes to learn the segmentation by extracting part-level equivariant features and factorizing the shape into canonical states. Through the iterative processing of canonicalization and then reconstruction using the factorized parameters, part decomposition can be learned by minimizing the reconstruction error on the input point cloud. PARIS liu2023paris, weng2024neural, and ArtGS liu2025building reconstruct articulated objects with part structure from a pair of multi-view RGB(-D) images that observe the object in two different states by leveraging differential rendering. By identifying the static and moving components through iterative optimization, the part-level geometry can be reconstructed by minimizing the photometric error between the input images and the rendered ones from the reconstructed surface. Following the trend, LEIA swaminathan2024leia proposes to reconstruct the interpolated object between different input states by leveraging a hypernetwork to modulate the network parameters of the reconstructed NeRF. They encode each state of the object into a latent code which can be linearly interpolated to generate the latent for the intermediate state, which is taken as the input to the hypernetwork to modulate the NeRF. This optimization process is supervised by the photometric loss to ensure the consistency of the reconstructed object and several regularization terms to enforce a smooth and continuous latent manifold. REACTO song2024reacto and RSRD kerr2024robot follows a similar idea to reconstruct the articulated objects from videos observing the object in motion. A photometric loss is used in these methods to learn the neural fields that represent the object with part-level structure.
Figure 11: Generative architecture based on diffusion model for unconditional articulated object synthesis from NAP lei2023nap. Figures reproduced from the original paper lei2023nap.
Generative models. Leveraging the recent advance of diffusion models and autoregressive models in 3D generation, a line of work has focused on training generative models to synthesize articulated objects with compositional part structures. NAP lei2023nap pioneers a novel representation that enables diffusion models to generate articulated objects with variational structures. As the diffusion model typically requires a fixed-size input, NAP represents the object as a set of parts padded to a fixed number, where each part is described by shape-motion attributes and its geometry encoded as a latent shape code, as shown in Figure 11. During generation, all parts are fully connected, and the model jointly synthesizes part attributes while reasoning about their dependencies to form a kinematic tree. The training follows a denoising diffusion process, where the model minimizes residual noise loss at each step to iteratively refine the generated structure. Once trained, new articulated object samples can be generated by progressive denoising from random noise. The denoiser is built on a graph convolutional network (GCN) architecture, which helps model hierarchical part relationships and capture kinematically plausible articulation structures. This framework showcases the effectiveness of diffusion models for articulated object generation, but the model shows limitations in user controllability over the generation process.
CAGE liu2024cage and SINGAPO liu2024singapo share a similar spirit to train diffusion models but focus on improving the controllability on the structure level and the user interaction in the generation process. These models tokenize part attributes and adopt a transformer-based architecture in the denoising network to capture part relationships via attention mechanisms. A key innovation is decoupling the part dependency graph from the output and using it as a guidance mechanism to enforce structural consistency throughout the generation process. CAGE proposes to use varying masking strategies on self-attention layers to enforce the graph conditional input effectively. Building on top of CAGE, SINGAPO introduces additional cross-attention blocks to inject spatial information into the part structure from an image feature map. This enables the model to generate multiple plausible articulated objects that are all consistent with an input image, improving the flexibility of image-conditioned generation. These methods demonstrate the potential of diffusion models for generating articulated objects with controllable structures and user-driven interactions. However, due to their part retrieval-based formulation, they lack the ability to synthesize diverse part geometries or ensure high compatibility between generated parts and the input image. Future work could explore integrating shape generation under a structural context, enabling more expressive and detailed synthesis to better align with input constraints and user intent.
Figure 12: Autoregressive model for mesh-based articulated object generation from MeshArt gao2024meshart. Figures reproduced from the original paper gao2024meshart.
Along with the generative approach built on diffusion models, another line of work explores generating articulated parts in an autoregressive manner. MeshArt gao2024meshart explores the articulated mesh generation in a part-by-part fashion. The process begins by generating a high-level articulation-aware object structure, which then guides the generation of individual part meshes. The key insight is to represent both structural information and part geometry as sequences of quantized triangle embeddings, making them well-suited for a unified autoregressive framework that captures articulation and fine-grained geometry jointly. This design allows the model to generate highly detailed meshes with realistic articulation, but it comes with certain limitations. The model is trained per object category, restricting its generalization across diverse articulated objects. Additionally, it assumes that articulation dependencies are limited to direct parent-child relationships (i.e., within one hop in the kinematic tree), which may not fully capture the complex hierarchical articulation structures present in more intricate objects. Similarly, ArtFormer su2024artformer adopts an autoregressive strategy to iteratively generate child parts for each known part. Each part is represented as a separate token with high-dimensional features, and feature chunks can be decoded into shape and articulation parameters. On the shape side, each part is represented by a bounding box and a latent code which is trained to be decoded into SDFs. A key advantage of this tokenization strategy is its simple alignment with text and image tokens, enabling the model to generate diverse articulated objects from user-provided textual or visual descriptions. This capability bridges text-to-3D synthesis with articulated object generation, offering a more interactive and controllable approach to structured 3D content creation.
Transfer learning methods. Trained on vast amounts of data, foundation models such as large language models (LLMs) and vision-language models (VLMs), along with general-purpose vision models like SAM kirillov2023sam and DUSt3R wang2024dust3r, have demonstrated significant potential in enhancing downstream tasks. These models serve as a base that can be adapted to a wide range of tasks with fine-tuning or prompting with a few examples. A recent line of work leverages these models to advance articulated object creation, utilizing their capabilities in understanding, segmenting, and reasoning about complex structures. liu2023few tackles the problem of few-shot articulated mesh generation by fine-tuning a convex deformation network pre-trained on part convex decomposition of the objects from similar categories. Real2Code mandi2024real2code proposes a framework to reconstruct articulated parts from multi-view images by fine-tuning SAM for segmenting articulated parts, and then lifting each part to a 3D point cloud using DUSt3R. Articulate-Anything le2024articulate proposes an actor-critic system to generate articulated part configurations using programs. Multiple VLM and LLM agents are prompted to specialize in different subtasks, such as object detection and reasoning from the visual input, part mesh retrieval and arrangement, rating the realism of the output, and providing specific feedback to correct the output. DragAPart li2024dragapart generates images of articulated objects constrained by a drag action by fine-tuning the stable diffusion model rombach2022high with paired drag-annotated images. Although trained on synthetic images, the model can generalize well to real images by leveraging the text-to-image prior in the pre-trained diffusion model. Articulate AnyMesh qiu2025anymesh proposes a pipeline to segment the input mesh by aggregating 2D segmentation results from multi-view images. The 2D articulated part segmentation is achieved by few-shoting the PartSlIP++ model zhou2023partslip using a few example masks for each category. It effectively leverages the advanced image segmentation models to assist the 3D segmentation task on surface meshes. For the occluded geometries, a refinement step is followed, which distills 3D geometry-texture priors from a pre-trained RichDreamer model qiu2024richdreamer using SDS loss poole2022dreamfusion.
These transfer learning based methods typically design a multi-stage pipeline that brings together the strengths of different foundation models to tackle each subproblem in the object creation process. Benefiting from powerful prior knowledge encoded in the base modules, these methods can potentially reduce the need for extensive annotated data to train a model from scratch, but only require a few examples to prompt the model or even used in zero-shot settings. However, the performance of these methods is highly dependent on the generalization capability and reliability of the base models. How robust the model can be adapted to the target task and how well it can generalize to unseen data are still open questions that need further exploration.
5 Articulation Modeling